OpenStudy (hpfan101):

How do you find the limit using L'Hospital's Rule? \[\lim_{x \rightarrow 0}\frac{ x-sinx }{ \tan^3x }\]

OpenStudy (misty1212):

hi take the derivative separately top and bottom, try again

OpenStudy (misty1212):

it won't work the first time, have to do it again

OpenStudy (misty1212):

first attempt gives \[\frac{1-\cos(x)}{3\tan^2(x)\sec^2(x)}\] still no good have to do it one more time

OpenStudy (hpfan101):

Oh okay, I'm going to show you what I got the second time to make sure I'm doing the derivative right

OpenStudy (hpfan101):

\[\frac{ sinx }{ 6tanx \times \sec^2x+3\tan^2x \times 2secx \times secx \times tanx }\]

OpenStudy (misty1212):

looks right,maybe there is a better way, because you still get 0/0

OpenStudy (hpfan101):

So would I take the derivative again? Or do I try to simplify?

OpenStudy (misty1212):

maybe we can do something else before that step

OpenStudy (hpfan101):

I was thinking about rewriting the 6tanx as (6)(sinx/cosx)? Not sure if that'll help

OpenStudy (misty1212):

ack i dont see a good way to change this to make it easier

OpenStudy (misty1212):

guess you have to go again

OpenStudy (zarkon):

\[\lim_{x\to0}\frac{1-\cos(x)}{3\tan^2(x)\sec^2(x)}=\lim_{x\to0}\frac{1-\cos(x)}{3\tan^2(x)}\lim_{x\to0}\frac{1}{\sec^2(x)}\]

OpenStudy (misty1212):

guess that is a good way!

OpenStudy (hpfan101):

Oh, so two different limits? Didn't know we can do that